OpenStudy (anonymous):

If cos Θ = negative two over five and tan Θ > 0, what is the value of sin Θ?

OpenStudy (anonymous):

Use \[ \cos^2\theta+\sin^2\theta = 1 \]

OpenStudy (anonymous):

That will give you the absolute value. The part about \(\tan\theta >0\) it just to let you know which quadrants it is in.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

I know it is not "Square Root of 21 over 5"

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle with cosine \(\frac{2}{5}\) you are missing the sine, which you do by finding the missing side via pythagoras

OpenStudy (anonymous):

It is -2/5

OpenStudy (anonymous):

The Cosine

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73???

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

did you find the length of the missing side?

OpenStudy (anonymous):

No, the cosine is -2/5

OpenStudy (anonymous):

You gave me an angle for 2/5

OpenStudy (anonymous):

fine you still need the length of the missing side i am not interested in whether it is positive or negative at the moment because i cannot draw a triangle with one side negative length we can worry about the sign of the answer later

OpenStudy (anonymous):

then the sin should equal The Square Root Of 29, negative or not.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

not even close

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

use pythagoras to find the missing side it is not \(\sqrt{29}\) for sure

OpenStudy (anonymous):

hint, it is a side, not the hypotenuse

OpenStudy (anonymous):

ohhh!

OpenStudy (anonymous):

what do you get for the side?

OpenStudy (anonymous):

the square root of 21

OpenStudy (anonymous):

right and that is not the sine, that is just the length of the side the sine is the opposite over the hypotenuse

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that makes your answer \[-\frac{\sqrt{21}}{5}\]

OpenStudy (anonymous):

i will let you figure out why it has to be negative

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw